I need help with solving 3x+y=-8 and -2x-y=6 by substitution.
Answer by mananth(16949)

3x+1y= -8 .............1
y=-8-3x

-2x-y= 6
Substitute y in (2)
-2x-1(-8-3x)=6
-2x+8+3x=6
-2x+3x= -2
x=-2
/ 1
x=-2
Plug the value of x in (1)
3x+y=-8
3*-2+y =-8
-6 + y = -8
1 y = -2
/ 1
y= -2
m.ananth@hotmail.ca